OpenStudy (anonymous):

Hexagon DEFGHI is translated on the coordinate plane below to create hexagon D'E'F'G'H'I': Hexagon DEFGHI and Hexagon D prime E prime F prime G prime H prime I prime on the coordinate plane with ordered pairs at D are 3, 5, at E 7, 5, at F 8, 2, at G 7, negative 1, at H 3, negative 1, at I 2, 2; at D prime negative 6, 2, at E prime negative 2, 2, at F prime negative 1, negative 1, at G prime negative 2, negative 4, at H prime negative 6, negative 4, at I prime negative 7, negative 1 Which rule represents the translation of hexagon DEFGHI to hexagon D'E'F'G'H'I'

OpenStudy (anonymous):

(x, y)(x - 9, y - 3) (x, y)(x - 3, y - 9) (x, y)(x + 3, y + 3) (x, y)(x + 9, y + 9)

OpenStudy (anonymous):

i think its a or b but always get the axis mixed yo some tips maybe?

OpenStudy (anonymous):

 i think it is a

OpenStudy (anonymous):

3-9=-6.5-3=2

OpenStudy (anonymous):

@surjithayer im so confused

OpenStudy (anonymous):

i have checked for D and D' D is (3,5) D' is (-6,2) for D x=3,for D' x=3-9=-6 for D ,y=5, for D' y=5-3=2

OpenStudy (anonymous):

hey are you there ?

OpenStudy (anonymous):

look at point F for example (8,2) it is translated to (-1,-1) so x'=x-9 and y'=y-3

OpenStudy (anonymous):

Thank you @Coolsector and @surjithayer

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

yw